A bar of mass M=2kg and length of .8 meters rotates about one of its endpoints with a constant angular velocity 88 rad/s

a) what is the kinetic energy of the system

b) what is the speed of the tip (given I=1/3 ML^2)
Answer #1

Here,

M = 2 Kg

L = 0.80 m

w = 88 rad/s

a) moment of inertia , I = 1/3 ML^2

kinetic energy of the system = 0.50 * I * w^2

kinetic energy of the system = 0.50 * 1/3 ML^2 * w^2

kinetic energy of the system = 0.50/3 * 2 * 0.80^2 * 88^2

kinetic energy of the system = 1652 J

b)

speed of the tip = L * w

speed of the tip = 88 * 0.80 m/s

speed of the tip = 70.4 m/s